# Just-In-Time Objectives: A General Approach for Specialized AI Interactions

### v1

@@ Abstract.
Large language models promise a broad set of functions, but when not given a specific objective, they default to milquetoast results such as drafting emails littered with cliches. We demonstrate that inferring the user’s in-the-moment objective, then rapidly optimizing for that singular objective, enables LLMs to produce tools, interfaces, and responses that are more responsive and desired. We contribute an architecture for automatically inducing just-in-time objectives by passively observing user behavior, then steering downstream AI systems through generation and evaluation against this objective. Inducing just-in-time objectives (e.g., “Clarify the abstract’s research contribution”) enables automatic generation of tools, e.g., those that critique a draft based on relevant HCI methodologies, anticipate related researchers’ reactions, or surface ambiguous terminology. In a series of experiments (N=14N={14}, N=205N={205}{}) on participants’ own tasks, JIT objectives enable LLM outputs that achieve 66–86% win rates over typical LLMs, and in-person use sessions (N=17) confirm that JIT objectives produce specialized tools unique to each participant.
human-AI interaction, JIT objectives, large language models
conference: ; 2025; preprintcopyright: noneccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Natural language interfaces
Figure 1. Just-in-time objectives are a method for producing AI objectives by inducing user goals on the fly from observations of the user and their task. We turn these objectives into first-class interactive objects that are visible, modifiable, and equipped to powerfully steer any number of downstream AI systems. By inducing possible objectives at interaction time, these objectives can be much more detailed and specialized to the particular task at hand compared to implicit, generic AI objectives.:
@@ 1. INTRODUCTION
In the course of writing this paper, we wondered what a large language model (LLM) might have to say about our introduction. The model offered helpful syntax edits (“replace technical phrases like hill-climb”), suggestions to aid readability (“some sentences could be broken up”), and thoughts on re-ordering content (“consider moving implementation specifics to a later section”). These suggestions all constitute useful advice, but we found them rather generic and unopinionated. Certainly a seasoned researcher would have more specific advice on our core argument.
@@
In this paper, we demonstrate that inferring the user’s in-the-moment objective, then rapidly optimizing the LLM’s behavior toward that singular objective, enables the LLM to generate outputs that are far more distinct and responsive—whether they be individual responses, extended drafts, or generated software tools. Instead of relying on an anticipatory objective that must be generic and opaque by default, we find that we can pursue an on-demand objective that is specific and malleable by default. User objectives over long stretches of time (e.g., a user’s writing philosophy) are complex and challenging to anticipate, which causes standard preemptive LLM objectives to fail. But when objectives are created on-demand based on small windows of observed user behavior, they become more tractable and accurate (e.g., a user’s aims for this phase of the Introduction). [^1] To make an analogy to calculus, even if a user’s objective is complex and curved over time, it can still be characterized as a simple straight line (a singular objective) over an infinitely small instant. We capture these instantaneous user goals as what we term just-in-time objectives: for example, “clarify the calculus example in the introduction,” “restructure the argument logic,” or “respond to Reviewer 2’s critique about sampling.” By inducing possible objectives at interaction time, these objectives can be much more detailed and specialized to the particular task at hand compared to implicit, generic AI objectives.
We contribute an architecture that automatically induces just-in-time objectives from interaction traces, then applies the just-in-time objective to steer AI behavior by shaping LLM generation and evaluation. This architecture induces strong shifts in model behavior: we demonstrate that applying this generate-and-evaluate algorithm to a just-in-time objective enables LLMs to rapidly construct tools that are specific and useful to the user in the moment. Recall that a standard LLM provided generic syntax and composition advice on an introduction like this one. Instead, applying an induced just-in-time objective of “explain the system clearly” might enable a system to automatically produce drafts that rework this paragraph’s logical flow to match that of similar introduction paragraphs from related HCI papers, or responses that simulate feedback from likely reviewers for the paper, or tools that identify where the narrative veers away from describing the system.
@@
To understand whether our just-in-time objectives are both correct and useful, we conduct a series of evaluations that assess the objectives in isolation and then incorporate them in increasingly complex LLM systems. We first run experiments on participant-submitted browser traces collected over three days (N=14N={14}{}) and find that across a variety of contexts in users’ daily lives, our just-in-time objectives are accurate and useful (acc: $M=2.04$, useful: $M=2.18$ on a -3 to +3 Likert scale, with 75.0% of ratings “2: Accurate/Useful” or higher). In addition, when applied to optimize LLM outputs that suggest feedback, areas of expertise, and tool designs, just-in-time objectives produce outputs strongly preferred, achieving win-rates between 71% and 86% over those of a baseline LLM. These results are corroborated by a large-scale experiment (N=205N={205}{}) across 410 participant-submitted input contexts, where just-in-time objectives are accurate, useful, and effective LLM optimizers (acc: $M=1.92$, useful: $M=2.06$, win-rates between 66% and 71% over baseline LLM). In fact, just-in-time objectives are selected as the most important objective for a task (as opposed to a custom-authored option) in 97.8% of cases. In hour-long study sessions applying Poppins to participants’ own writing tasks (N=17N={17}{}), we find that generated tools constructed using just-in-time objectives are more relevant and useful than those served by a baseline LLM chat tool. Poppins produces a broad range of tailored tools including a “Cultural Perspective Highlighter” for a scholarship personal essay, “Neural Architecture Search (NAS) Explorer” for a research project on microcontrollers, “Technical Protocol Generator” for a bioengineering research proposal, and “Character Emotion Tracker” for a science fiction short story draft.
@@
  • An architecture for just-in-time objectives: AI objectives that are automatically induced in response to a user’s need. Our method centers around a goal-based generation and evaluation loop to optimize AI output without supervision.
  • The Poppins system, which instantiates our vision of just-in-time objectives in the domain of generative user interfaces by translating observed user contexts to functional interactive tools. With Poppins, just-in-time objectives generate tools by inducing goal-aligned design specifications.
  • Evaluations that validate the accuracy and utility of just-in-time objectives and outputs of the AI systems they optimize, such as expert feedback and generated software tools.
@@ 2. RELATED WORK
Our work builds on evolving explorations of human-LLM interaction, as well as an emerging literature that names worrying individual- and population-level failure modes. Our approach works towards a goal of interactive systems that specialize on the fly—echoing both classic literature on adaptive interfaces and recent work on dynamic user interface generation.
@@ 2.1. Failure Modes of Human-LLM Interaction: Undirected Prompting and Generic Outputs
Another strong motivator of our work is the problem of generic LLMs. Emerging research presents evidence of the homogenizing impacts of generic LLM outputs for both individual-level (Padmakumar and He, 2024; Lee et al., 2025; Kumar et al., 2025; Chakrabarty et al., 2024) and population-level creativity (Anderson et al., 2024; Wu et al., 2024b; Doshi and Hauser, 2024; Wenger and Kenett, 2025; Messeri and Crockett, 2024). From LLMs, we observe an uncanny valley of model behaviors that appear divergent, but are not fully random, and yet effortful to meaningfully steer (Ugander and Epstein, 2024). Worryingly, at a population level, LLMs appear to promote monocultures, steering users towards homogeneous, convergent thinking even when individual outputs appear creative (Anderson et al., 2024; Doshi and Hauser, 2024). One response is that we need pluralistic modeling approaches that better represent the diversity of the population (Feng et al., 2024b, a) and account for pluralistic values (Sorensen et al., 2024a, b). Another response is that we need mechanisms for greater and more sustainable participation in the development and maintenance of LLMs (Suresh et al., 2024; Huang et al., 2024). We share a concern about the risks of current LLM tools promoting monocultures, and our work explores whether even generic models can be effectively coaxed into specific, divergent variants with just-in-time objectives. We posit that if we can grant users on-demand access to specific, divergent model behaviors, they may be able to break away from generic model outputs and actualize their unique creative voice.
@@ 2.2. Adaptive UIs & User Modeling: Building on Past Lessons
While in theory, adaptive interfaces seem unilaterally better than static interfaces, in practice, adaptive UIs face common pitfalls around unpredictability and loss of user control (Findlater and Gajos, 2009). With recent developments in LLMs, a number of recent works expand the vision of adaptive interfaces by supporting dynamic widgets and on-the-fly code execution in contexts such as data visualization and computational notebook tutorials (Vaithilingam et al., 2024; Cheng et al., 2024; Dibia, 2023). While our work does not rid itself entirely of unpredictability due to the use of black-box LLMs, we argue that conjuring objective-aligned tools from a vast set of options (and making this objective visible) elides many of the traditional issues of adaptive UIs, which stemmed from a finite set of interactive functions moving around in unpredictable ways.
@@ 2.3. Dynamic UI Generation: Direction-Setting With the Right Optimizer
We share an excitement about this research direction, while also noting that achieving a “dynamic” interface is not inherently valuable: interfaces must adapt along the axes that users find valuable. Recent work on UI generation demonstrates the importance of feedback to steer UI generation towards high-quality designs. Large-scale data such as eye-tracking data (Jiang et al., 2023) and repositories of existing UIs (Wu et al., 2025, 2024c) can generate reliable feedback, but such data is costly to gather. We draw inspiration from self-improving algorithms for LLM pipelines (Madaan et al., 2023) that leverage lightweight metrics rather than massive datasets (Khattab et al., 2024; Zhu et al., 2025), but our work explores whether these metrics could be directly induced from passive observations of user behavior. Just-in-time objectives provide a means for users to express what axes they find valuable and guide UI generation—without requiring manual customization.
@@ 3. JUST-IN-TIME OBJECTIVES
With a critical role in shaping model behavior from conversational chatbots to code assistants to writing tools, large language model objectives are often fixed far in advance of the scenarios they will encounter. In-context prompting can overcome this issue, but users are rarely detailed enough in their request to cover all required decisions (Shaikh et al., 2024). In this section, we describe an architecture that shifts the work of objective-making to the site of the user, at the specific moment of need. Here, we describe a lightweight architecture to achieve this goal when developing LLM applications. Users are much more complex and challenging to model than is implicitly assumed by standard approaches that make monolithic, static assumptions about an individual’s objective—however, we can can sidestep this thorny challenge by waiting until interaction time to ground this inference.
In the following sections, we walk through the technical architecture of just-in-time (JIT) objectives that can be implemented into any interactive system. We demonstrate how crisp objectives in-the-moment can shape AI applications, taking examples of common system components responsible for generation and evaluation. We then illustrate how user-facing systems can be reimplemented with just-in-time objectives by introducing the Poppins system, which incorporates just-in-time objectives to drive customized UI generation.
@@ 3.1. Architecture
In this section, we work backwards by first demonstrating (1) how such an objective can be minimally applied to common components of LLM systems to steer their behavior, and then discuss (2) a process for creating just-in-time objectives by observing user behavior.
Figure 2. Applying Just-In-Time Objectives. The core of our architecture is a just-in-time objective that can add on to LLM system components, such as those responsible for generation and evaluation. Here, we show an example of dual generation (gen_objective) and evaluation (eval_objective) steps that optimize model behavior. The generation step produces candidates of the specified output format based on the user context and goal, and the evaluation step assesses these candidates with respect to the goal. :
@@ 3.1.1. Generate: Applying JIT Objectives to Optimize Generation
Mary enables objective induction based on editing patterns in her paper and appends the result to her existing prompt. Now, she gets more varied and interesting suggestions. The system induced an objective of “Strengthen the narrative argument,” which steered the system to provide advice on how to move beyond sequentially describing the system components and instead present a narrative walkthrough of a user applying the system.
gen_objective can apply to any generation call, so objectives can also be applied to generators that revise existing content, initiate tool calls, or perform actions on behalf of the user.
Figure 3. The JSON specification for just-in-time objectives includes a name, 1-2 sentence description, and weight indicating the estimated importance of the objective (1-10 scale).:

### v2

@@ Abstract.
Large language models promise a broad set of functions, but when not given a specific objective, they default to generic results. We demonstrate that inferring the user’s in-the-moment objective, then rapidly optimizing for that singular objective, enables LLMs to produce specialized tools, interfaces, and responses. Our work introduces just-in-time objectives, which model a user’s goals to specialize LLM systems on the fly. We contribute an architecture for automatically inducing such objectives by passively observing user behavior, then steering downstream AI systems through generation and evaluation against this objective. Inducing just-in-time objectives (e.g., “Clarify the abstract’s research contribution”) enables automatic generation of tools, e.g., those that critique a draft based on relevant HCI methodologies, anticipate related researchers’ reactions, or surface ambiguous terminology. In a series of experiments on participants’ own tasks, JIT objectives enable LLM outputs that achieve 66–86% win rates over typical LLMs. In-person use sessions confirm that JIT objectives produce specialized tools that are unique to each participant and are rated as significantly higher quality than a standard LLM chat tool.
human-AI interaction, just-in-time objectives, JIT objectives, large language models
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790713isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Natural language interfaces
Figure 1. Just-in-time objectives are a method for producing AI objectives by inducing user goals on the fly from observations of the user and their task. We turn these objectives into first-class interactive objects that are visible, modifiable, and equipped to powerfully steer any number of downstream AI systems. By inducing possible objectives at interaction time, these objectives can be much more detailed and specialized to the particular task at hand compared to implicit, generic AI objectives.
A conceptual diagram showing the flow from user context observation to objective induction to specialized AI system outputs. The left side shows various windows of a user’s workspace that have been passively observed. The center shows the just-in-time objectives that feed into generation and evaluation operations. The right side displays various AI-generated outputs including tools and interfaces that are more detailed and specialized compared to generic large language model outputs.
@@ 1. INTRODUCTION
In the course of writing this paper, we wondered what a large language model (LLM) might have to say about our introduction. The model offered helpful syntax edits (“replace technical phrases like hill-climb”), suggestions to aid readability (“some sentences could be broken up”), and thoughts on re-ordering content (“consider moving implementation specifics to a later section”). These suggestions all constitute useful advice, but we found them rather milquetoast and unopinionated. Certainly a seasoned researcher would have more specific advice on our core argument.
@@
In this paper, we demonstrate that inferring the user’s in-the-moment objective, then rapidly optimizing the LLM’s behavior toward that singular objective, enables the LLM to generate outputs that are far more distinct and responsive: whether they be individual responses, extended drafts, or generated software tools. Instead of relying on an anticipatory objective that must be generic and opaque by default, we find that we can pursue an on-demand objective that is specific and malleable by default. User objectives over long stretches of time (e.g., a user’s writing philosophy) are complex and challenging to anticipate, which causes standard preemptive LLM objectives to fail. But when objectives are created on-demand based on small windows of observed user behavior, they become more tractable and accurate (e.g., a user’s aims for this phase of the Introduction). [^1] To make an analogy to calculus, even if a user’s objective is complex and curved over time, it can still be characterized as a simple straight line (a singular objective) over an infinitely small instant. We capture these instantaneous user goals as what we term just-in-time objectives: for example, “clarify the calculus example in the introduction,” “restructure the argument logic,” or “respond to Reviewer 2’s critique about sampling.” By inducing possible objectives at interaction time, these objectives can be much more detailed and specialized to the particular task at hand compared to implicit, generic AI objectives.
We contribute a generalizable architecture for user-specific LLM specialization that automatically induces just-in-time (JIT) objectives from interaction traces, then applies the JIT objective to steer AI behavior by shaping LLM generation and evaluation. This architecture induces strong shifts in model behavior across task domains: we demonstrate that applying a JIT objective in a generate-then-rank architecture enables LLMs to rapidly construct tools that are specific and useful to the user in the moment. Recall that a standard LLM provided generic syntax and composition advice on an introduction like this one. Instead, applying an induced just-in-time objective of “explain the system clearly in this CHI paper introduction” enables a system to automatically produce drafts that rework this paragraph’s logical flow to match that of similar introduction paragraphs from related CHI papers, or responses that simulate feedback from likely CHI reviewers, or tools that identify where the narrative veers away from describing the system.
@@
To understand whether our just-in-time objectives are both accurate and useful, we conduct a series of evaluations that assess the objectives in isolation and then incorporate them in increasingly complex LLM systems. We first run experiments on participant-submitted browser traces collected over three days (N=14N={14}{}) and find that across a variety of contexts in users’ daily lives, our just-in-time objectives are accurate and useful (acc: $M=2.04$, useful: $M=2.18$ on a -3 to +3 Likert scale, with 75.0% of ratings “2: Accurate/Useful” or higher). In addition, when applied to optimize LLM outputs that suggest feedback, areas of expertise, and tool designs, just-in-time objectives produce strongly-preferred outputs, achieving win-rates between 71% and 86% over those of a baseline LLM. These results are corroborated by a large-scale experiment (N=205N={205}{}) across 410 participant-submitted input contexts, where just-in-time objectives are accurate, useful, and effective LLM optimizers (acc: $M=1.92$, useful: $M=2.06$, win-rates between 66% and 71% over baseline LLM). In hour-long study sessions applying Poppins to participants’ own writing tasks (N=17N={17}{}), we find that generated tools constructed using just-in-time objectives are relevant and useful without requiring manual prompting. Poppins outputs yield overall quality ratings significantly higher than a standard LLM chat tool ($p<.05$). Poppins produces a broad range of tailored tools including a “Cultural Perspective Highlighter” for a scholarship personal essay, “Neural Architecture Search (NAS) Explorer” for a research project on microcontrollers, “Technical Protocol Generator” for a bioengineering research proposal, and “Character Emotion Tracker” for a science fiction short story draft.
@@
  • A generalizable architecture for just-in-time objectives: a model of a user’s in-the-moment goals that is automatically induced to achieve user-specific LLM specialization. Our method centers around an objective-based generation and evaluation loop to optimize AI output without supervision.
  • The Poppins [^2] Code available at: https://github.com/michelle123lam/poppins system, which instantiates our vision of just-in-time objectives in the domain of generative user interfaces by translating observed user contexts to functional interactive tools. With Poppins, just-in-time objectives generate tools by inducing objective-aligned design specifications.
  • Evaluations that validate the accuracy and utility of just-in-time objectives and outputs of the AI systems they optimize, such as expert feedback and generated software tools in Poppins.
@@ 2. RELATED WORK
Our work builds on evolving explorations of human-LLM interaction, as well as an emerging literature that names worrying individual- and population-level failure modes. Our approach works towards a goal of interactive systems that specialize on the fly, echoing classic literature on adaptive interfaces and recent work on user interface generation.
@@ 2.1. Failure Modes of Human-LLM Interaction: Undirected Prompting and Generic Outputs
Another strong motivator of our work is the problem of generic LLMs. Emerging research presents evidence of the homogenizing impacts of generic LLM outputs for both individual-level (Padmakumar and He, 2024; Lee et al., 2025; Kumar et al., 2025; Chakrabarty et al., 2024) and population-level creativity (Anderson et al., 2024; Wu et al., 2024a; Doshi and Hauser, 2024; Wenger and Kenett, 2025; Messeri and Crockett, 2024). From LLMs, we observe an uncanny valley of model behaviors that appear divergent, but are not fully random, and yet effortful to meaningfully steer (Ugander and Epstein, 2024). Worryingly, at a population level, LLMs appear to promote monocultures, steering users towards homogeneous, convergent thinking even when individual outputs appear creative (Anderson et al., 2024; Doshi and Hauser, 2024). One response is that we need pluralistic modeling approaches that better represent the diversity of the population (Feng et al., 2024b, a) and account for pluralistic values (Sorensen et al., 2024a, b). Another response is that we need mechanisms for greater and more sustainable participation in the development and maintenance of LLMs (Suresh et al., 2024; Huang et al., 2024). We share a concern about the risks of current LLM tools promoting monocultures, and our work explores whether even generic models can be effectively coaxed into specific, divergent variants with just-in-time objectives. In contrast with approaches that specialize LLM behavior with domain- or task-specific context (Wadhwa et al., 2025; Lin et al., 2024), our approach aims for user-specific specialization that can generalize across tasks and domains. We posit that if we can grant users on-demand access to specific, divergent model behaviors, they may be able to break away from generic model outputs and actualize their unique creative voice.
@@ 2.2. Adaptive UIs & User Modeling: Building on Past Lessons
While in theory, adaptive interfaces seem unilaterally better than static interfaces, in practice, adaptive UIs face common pitfalls around unpredictability and loss of user control (Findlater and Gajos, 2009). With developments in LLMs, a number of recent works expand the vision of adaptive interfaces by supporting dynamic widgets and on-the-fly code execution in contexts such as data visualization and computational notebook tutorials (Vaithilingam et al., 2024; Cheng et al., 2024; Dibia, 2023). While our work does not rid itself entirely of unpredictability due to the use of black-box LLMs, we argue that conjuring objective-aligned tools from a vast set of options (and making this objective visible) elides many of the traditional issues of adaptive UIs, which stemmed from a finite set of interactive functions moving around in unpredictable ways.
@@ 2.3. Dynamic UI Generation: Direction-Setting With the Right Optimizer
We share an excitement about this research direction, while also noting that achieving a “dynamic” interface is not inherently valuable: interfaces must adapt along the axes that users find valuable. Recent work on UI generation demonstrates the importance of feedback to steer UI generation towards high-quality designs. Large-scale data such as eye-tracking data (Jiang et al., 2023) and repositories of existing UIs (Wu et al., 2025, 2024b) can generate reliable feedback, but such data is costly to gather. We draw inspiration from self-improving algorithms for LLM pipelines (Madaan et al., 2023) that leverage lightweight metrics rather than massive datasets (Khattab et al., 2024; Zhu et al., 2025), but our work explores whether these metrics could be directly induced from passive observations of user behavior. Given such user-specific metrics, it is possible to build on the success of LLM generate-then-rank architectures that can scale up LLM performance at inference time with generators that produce a large number of samples and evaluators (or verifiers) that select the best output candidate (Brown et al., 2024; Song et al., 2025; Snell et al., 2024; Saad-Falcon et al., 2025). Just-in-time objectives provide a means for users to express what axes they find valuable and guide UI generation, without requiring manual customization.
@@ 3. JUST-IN-TIME OBJECTIVES
With a critical role in shaping model behavior from conversational chatbots to code assistants to writing tools, large language model objectives are often fixed far in advance of the scenarios they will encounter. In-context prompting can overcome this issue, but users are rarely detailed enough in their request to cover all required decisions (Shaikh et al., 2024). In this section, we describe an architecture that shifts the work of objective-making to the site of the user, at the specific moment of need.
Our work contributes a generalizable architecture to achieve user-specific LLM specialization:
  • Our primary goal is user-specific LLM specialization: LLM output that best aligns with a particular user’s needs. Domain- or task-specific needs can be gleaned from the web, expert-authored checklists, or model self-reflection as in prior work (Madaan et al., 2023; Wadhwa et al., 2025; Lin et al., 2024). By contrast, user-specific needs depend on not just the domain (e.g., writing) or task (e.g., outlining), but factors such as the user’s specific task instance (e.g., the content, audience, and current progress of their outline), task environment (e.g., current application or tool, visible portion of outline), and personal preferences. All of these factors strongly shape a user’s current objective and can only be gathered from firsthand data about their specific interaction context.
  • To achieve this goal, we contribute a generalizable architecture that functions across task domains. Ideas of user-specific specialization exist in individual domains (e.g., writing assistance, persona generation, web search (Benharrak et al., 2024; D’Arcy et al., 2024; Choi et al., 2025; Baek et al., 2024; Salemi et al., 2024)), but are implemented with domain-specific assumptions and are not designed to transfer across domains. Our architecture can consistently and automatically apply user-specific specialization without domain-specific assumptions or manual steering from users or developers.
In the following sections, we walk through the technical architecture of just-in-time (JIT) objectives that can be implemented into any interactive system. We demonstrate how crisp objectives in-the-moment can shape AI applications, taking examples of common system components responsible for generation, evaluation, and iterative improvement. We then illustrate how user-facing systems can be reimplemented with just-in-time objectives by introducing the Poppins system, which incorporates just-in-time objectives to drive customized UI generation.
@@ 3.1. Architecture
In this section, we work backwards by first demonstrating (1) how such an objective can be minimally applied to common components of LLM systems to steer their behavior, and then discuss (2) a process for creating just-in-time objectives by observing user behavior. In particular, our architecture builds on the existing design pattern of generators and evaluators in AI systems. These are also known as generators and verifiers (Saad-Falcon et al., 2025; Brown et al., 2024; Song et al., 2025; Snell et al., 2024) or generate-then-rank methods  (Zhang et al., 2023; Xu et al., 2020), with analogues to actor-critic (Konda and Tsitsiklis, 1999) and generator-discriminator architectures (Goodfellow et al., 2014) for prior AI systems. We design our system so that induced objectives can slot into existing generators and evaluators in LLM systems.
Figure 2. Applying Just-In-Time Objectives. The core of our architecture is a just-in-time objective that can add on to LLM system components, such as those responsible for generation and evaluation. Here, we show an example of dual generation (gen_objective) and evaluation (eval_objective) steps that optimize model behavior. The generation step produces candidates of the specified output format based on the user context and objective, and the evaluation step assesses these candidates with respect to the objective.
Architecture diagram showing how just-in-time objectives enhance generation and evaluation components in language model systems. A flowchart depicts the core architecture with three main components: (1) inputs of User context and Output format, (2) a core operation of Generation and Evaluation, and (3) Output resulting from this process.
@@ 3.1.1. Generate: Applying JIT Objectives to Optimize Generation
Mary enables objective induction based on editing patterns in her paper and appends the result to her existing prompt. Now, she gets more varied and interesting suggestions. The system induced an objective of “Strengthen the narrative argument,” which steered the system to provide advice on how to move beyond sequentially describing system components and instead present a narrative user walkthrough of the system.
The gen_objective operator can apply to any generation call, so objectives can also be applied to generators that revise existing content, initiate tool calls, or perform actions on behalf of the user.
Figure 3. The JSON specification for just-in-time objectives includes a name, 1-2 sentence description, and weight indicating the estimated importance of the objective on a 1-10 scale.
JSON specification format for just-in-time objectives containing name, description, and weight fields.